OpenStudy (anonymous):

An ellipse with foci at (-6, -3) and (0, -3); major axis length of 10

OpenStudy (anonymous):

okay, what about it? you need to ask a question to get an answer.

OpenStudy (anonymous):

write equation about that ?

OpenStudy (amistre64):

(y+3)^2 (x+3)^2 ------- + -------- = 1 a^2 b^2

OpenStudy (amistre64):

c = 3; a = 5

OpenStudy (amistre64):

a^2 = 25 b^2 = a^2 - c^2 b^2 = 25 - 9 = 16